OpenStudy (anonymous):

Nathan bought a pair of boots for $79.95 and 2 shirts for $19.99 each. The sales tax is 7.5%. How much did Nathan pay for all the items, including tax? A $99.94 B $119.93 C $128.92 D $147.95

OpenStudy (laurenmyers1098):

You asked this in the English section, you could probably get more help by asking in the Mathematics section! Anyway, the answer is C.

OpenStudy (laurenmyers1098):

19.99 x 2 = 39.98 39.98 + 79.95 = 119.93 7.5% of 119.93 = 8.99475 8.99475 + 119.93 = 128.92475 = 128.92

OpenStudy (texaschic101):

(79.95 + 2(19.99)) * 0.075 = your answer

OpenStudy (anonymous):

THanks!